# Arithmetic Bohr radius and Local Banach space theory

### v1

@@ 5. Proof of the main results
$\left(1+\frac{n}{m}\right)^{-\frac{(s-1)mt-s+t}{mst}}\geq\frac{1}{2}\,\min\left\{1,\left(\frac{n}{m}\right)^{-\frac{(s-1)m-(\frac{s}{t}-1)}{ms}}\right\}=\frac{1}{2}\,\min\left\{1,\frac{m^{\frac{s-1}{s}}n^{\frac{(\frac{s}{t}-1)}{sm}}}{n^{\frac{s-1}{s}}m^{\frac{(\frac{s}{t}-1)}{sm}}}\right\}.$ (5.20)
It is not difficult to see that the function $\psi:(0,\infty)\rightarrow\mathbb{R}$ defined by $\psi(x)=x^{\frac{s-1}{s}}n^{\frac{(\frac{s}{t}-1)}{sm}}$, attains a strict minimum at $x=\log\,n$ with the minimum value $e^{\frac{(\frac{s}{t}-1)}{s}}\,(\log\,n)^{\frac{s-1}{s}}$. Finally, this fact together with Lemma 4.2 (1), (5.19), and (5.20), yields
$R_{\lambda}(B_{Z},p,U)\geq D\,\left(\frac{\lambda^{p}-\left\lVert U\right\rVert^{p}}{2\lambda^{p}-\left\lVert U\right\rVert^{p}}\right)^{\frac{1}{p}}\,\left(\frac{\log\,n}{n}\right)^{1-\frac{1}{s}}$ (5.21)
@@
and hence we obtain $(R_{\lambda}(\mathbb{D},p,U)/n)\leq AP_{\lambda}(B_{Z},p,U)$. Conversely, we prove that
@@
$\sum_{m=0}^{\infty}\sum_{|\alpha|=m}(\left\lVert U(a_{\alpha})\right\rVert^{p}_{Y}+\left\lVert U(b_{\alpha})\right\rVert^{p}_{Y})r^{p\alpha}\leq\lambda^{p}\left\lVert f\right\rVert^{p}_{\Omega,X}.$ (5.22)
@@
Let $\tilde{f}:\mathbb{D}\rightarrow X\in\mathcal{PH}(\mathbb{D},X)$ be a bounded holomorphic map of the form $\tilde{f}(\xi)=\sum_{m=0}^{\infty}c_{m}\xi^{m}+\sum_{m=1}^{\infty}d^{*}_{m}\overline{\xi}^{m}$ for $\xi\in\mathbb{D}$. We now consider the function $\phi:B_{Z}\rightarrow\mathbb{C}$ defined by
$\phi(z)=\frac{z_{1}+\cdots+z_{n}}{t},\quad z\in B_{Z},$ (5.23)
@@
Acknowledgment. The author gratefully acknowledges the support received during the initial stages of this work while serving as a Senior Project Associate at IIT Bombay, through the Core Research Grant (CRG), sanctioned to Professor Sourav Pal by the Science and Engineering Research Board (SERB), Government of India. The author is currently supported by the J. C. Bose Fellowship of Professor B. V. Rajarama Bhat, ISI Bangalore, whose support is also sincerely acknowledged.

### v2

@@ 5. Proof of the main results
(5.20) $\left(1+\frac{n}{m}\right)^{-\frac{(s-1)mt-s+t}{mst}}\geq\frac{1}{2}\,\min\left\{1,\left(\frac{n}{m}\right)^{-\frac{(s-1)m-(\frac{s}{t}-1)}{ms}}\right\}=\frac{1}{2}\,\min\left\{1,\frac{m^{\frac{s-1}{s}}n^{\frac{(\frac{s}{t}-1)}{sm}}}{n^{\frac{s-1}{s}}m^{\frac{(\frac{s}{t}-1)}{sm}}}\right\}.$
It is not difficult to see that the function $\psi:(0,\infty)\rightarrow\mathbb{R}$ defined by
$\psi(x)=x^{\frac{s-1}{s}}n^{\frac{(\frac{s}{t}-1)}{sm}},$
attains a strict minimum at $x=\log\,n$ with the minimum value
$e^{\frac{(\frac{s}{t}-1)}{s}}\,(\log\,n)^{\frac{s-1}{s}}.$
Finally, this fact together with Lemma 4.2 (1), (5.19), and (5.20), yields
(5.21) $R_{\lambda}(B_{Z},p,U)\geq D\,\left(\frac{\lambda^{p}-\left\lVert U\right\rVert^{p}}{2\lambda^{p}-\left\lVert U\right\rVert^{p}}\right)^{\frac{1}{p}}\,\left(\frac{\log\,n}{n}\right)^{1-\frac{1}{s}}$
@@
and hence we obtain
$\frac{R_{\lambda}(\mathbb{D},p,U)}{n}\leq AP_{\lambda}(B_{Z},p,U).$
Conversely, we prove that
@@
(5.22) $\sum_{m=0}^{\infty}\sum_{|\alpha|=m}(\left\lVert U(a_{\alpha})\right\rVert^{p}_{Y}+\left\lVert U(b_{\alpha})\right\rVert^{p}_{Y})r^{p\alpha}\leq\lambda^{p}\left\lVert f\right\rVert^{p}_{\Omega,X}.$
@@
Let $\tilde{f}:\mathbb{D}\rightarrow X\in\mathcal{PH}(\mathbb{D},X)$ be a bounded holomorphic map of the form
$\tilde{f}(\xi)=\sum_{m=0}^{\infty}c_{m}\xi^{m}+\sum_{m=1}^{\infty}d^{*}_{m}\overline{\xi}^{m}$
for $\xi\in\mathbb{D}$. We now consider the function $\phi:B_{Z}\rightarrow\mathbb{C}$ defined by
(5.23) $\phi(z)=\frac{z_{1}+\cdots+z_{n}}{t},\quad z\in B_{Z},$
@@
Proof of Lemma 1.1. Since $U$ is the identity operator on $\mathbb{C}$, we consider the complex valued harmonic functions on $\mathbb{D}$. Let $f:\mathbb{D}\rightarrow\mathbb{C}$ be a bounded harmonic function of the form
$f(z)=\sum_{n=0}^{\infty}a_{n}\,z^{n}+\overline{\sum_{n=1}^{\infty}b_{n}\,z^{n}},\,\,\,z\in\mathbb{D}.$
Then, by [1, Lemma 4], we have
$|a_{n}|+|b_{n}|\leq\frac{4}{\pi}\,\left\lVert f\right\rVert_{\mathbb{D},\mathbb{C}}\,\,\,\,\,\mbox{for}\,\,\,n\geq 1.$
By virtue of this estimate, for $|z|=r<1$, we obtain
$|a_{0}|+\sum_{n=1}^{\infty}(|a_{n}|+|b_{n}|)r^{n}$ $\leq|a_{0}|+\frac{4}{\pi}\,\frac{r}{1-r}\,\left\lVert f\right\rVert_{\mathbb{D},\mathbb{C}}$
$\leq\left\lVert f\right\rVert_{\mathbb{D},\mathbb{C}}+\frac{4}{\pi}\,\frac{r}{1-r}\,\left\lVert f\right\rVert_{\mathbb{D},\mathbb{C}}$
$=\left(1+\frac{4}{\pi}\,\frac{r}{1-r}\right)\,\left\lVert f\right\rVert_{\mathbb{D},\mathbb{C}}\leq\lambda\,\left\lVert f\right\rVert_{\mathbb{D},\mathbb{C}},$
provided
$r\leq\frac{(\lambda-1)\pi}{4+(\lambda-1)\pi}.$
Hence, the desired conclusion follows. The proof is now complete. ∎
@@
Acknowledgment. The author gratefully acknowledges the support received during the initial stages of this work while serving as a Senior Project Associate at IIT Bombay, through the Core Research Grant (CRG), sanctioned to Professor Sourav Pal by the Science and Engineering Research Board (SERB), Government of India. The author also sincerely acknowledges Professor B. V. Rajarama Bhat, ISI Bangalore, for support through the J. C. Bose Fellowship during the final stages of this manuscript. The author is currently supported by an NBHM Postdoctoral Fellowship from the Department of Atomic Energy (DAE), Government of India.